The width of a rectangular poster is one-half of its length. The area of the poster is 128 sq. ft. Click on the equation that will determine the length.
Let L = length.

Answer by rfer(16322)

x*2x=128
2x^2=128
x^2=64
x=8 ft
2x=16 ft